# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ B Comparative Details on 216 Instances of SET-1
brock200_2_A 200 4916 4960 1298 8390 1298 93 1298 $<$1 1298 $<$1 1298 $<$1
brock400_2_A 400 29711 30075 1103 10800 1123 355 1123 $<$1 1123 $<$1 1123 $<$1
C125.9_A 125 3500 3463 627 5 627 $<$1 627 $<$1 627 $<$1 627 $<$1
C250.9_A 250 14017 13967 817 266 817 1 817 $<$1 817 $<$1 817 $<$1
gen200_p0.9_55_A 200 8908 9002 785 68 785 $<$1 785 $<$1 785 $<$1 785 $<$1
gen400_p0.9_75_A 400 35823 35997 895 5699 895 29 895 $<$1 895 $<$1 895 $<$1
hamming8-4_A 256 10329 10535 1301 7608 1301 49 1301 $<$1 1301 $<$1 1301 $<$1
keller4_A 171 4738 4697 1118 153 1118 2 1118 $<$1 1118 $<$1 1118 $<$1
MANN_a27_A 378 35345 35206 812 1217 812 6 812 $<$1 812 $<$1 812 $<$1
p_hat300-1_A 300 5505 5428 2568 10800 3026 10800 3129 $<$1 3129 $<$1 3129 $<$1
p_hat300-2_A 300 11051 10877 2063 10800 2477 10800 2477 $<$1 2477 $<$1 2477 $<$1
p_hat300-3_A 300 16820 16570 1029 3748 1029 38 1029 $<$1 1029 $<$1 1029 $<$1
brock200_2_B 200 4916 4960 1224 9589 1224 125 1224 $<$1 1224 $<$1 1224 $<$1
brock400_2_B 400 29711 30075 1010 10800 1035 445 1035 $<$1 1035 $<$1 1035 $<$1
C125.9_B 125 3500 3463 582 5 582 $<$1 582 $<$1 582 $<$1 582 $<$1
C250.9_B 250 14017 13967 744 292 744 2 744 $<$1 744 $<$1 744 $<$1
gen200_p0.9_55_B 200 8908 9002 716 72 716 $<$1 716 $<$1 716 $<$1 716 $<$1
gen400_p0.9_75_B 400 35823 35997 805 6602 805 34 805 $<$1 805 $<$1 805 $<$1
hamming8-4_B 256 10329 10535 1255 7780 1255 62 1255 $<$1 1255 $<$1 1255 $<$1
keller4_B 171 4738 4697 1094 156 1094 2 1094 $<$1 1094 $<$1 1094 $<$1
MANN_a27_B 378 35345 35206 707 1417 707 7 707 1 707 $<$1 707 $<$1
p_hat300-1_B 300 5505 5428 2492 10800 2950 10800 3023 $<$1 3023 $<$1 3023 $<$1
p_hat300-2_B 300 11051 10877 1979 10800 2405 10800 2405 $<$1 2405 $<$1 2405 $<$1
p_hat300-3_B 300 16820 16570 967 4067 967 46 967 $<$1 967 $<$1 967 $<$1
brock200_2_C 200 4916 4960 1101 10800 1101 204 1101 $<$1 1101 $<$1 1101 $<$1
brock400_2_C 400 29711 30075 861 10800 892 691 892 $<$1 892 $<$1 892 $<$1
C125.9_C 125 3500 3463 506 6 506 $<$1 506 $<$1 506 $<$1 506 $<$1
C250.9_C 250 14017 13967 623 334 623 2 623 $<$1 623 $<$1 623 $<$1
gen200_p0.9_55_C 200 8908 9002 597 91 597 $<$1 597 $<$1 597 $<$1 597 $<$1
gen400_p0.9_75_C 400 35823 35997 651 7858 651 50 651 $<$1 651 $<$1 651 $<$1
hamming8-4_C 256 10329 10535 1184 7447 1184 85 1184 $<$1 1184 $<$1 1184 $<$1
keller4_C 171 4738 4697 1049 145 1049 3 1049 $<$1 1049 $<$1 1049 $<$1
MANN_a27_C 378 35345 35206 552 1849 552 10 552 1 552 $<$1 552 $<$1
p_hat300-1_C 300 5505 5428 2303 10800 2818 10800 2897 $<$1 2897 $<$1 2897 $<$1
p_hat300-2_C 300 11051 10877 1852 10800 2263 10800 2263 $<$1 2263 $<$1 2263 $<$1
p_hat300-3_C 300 16820 16570 851 4709 851 69 851 $<$1 851 $<$1 851 $<$1
@@
brock200_2_A 200 4916 4960 2034 10800 2034 677 2034 $<$1 2034 $<$1 2034 $<$1
brock400_2_A 400 29711 30075 1628 10800 1630 1323 1630 11 1630 $<$1 1630 $<$1
C125.9_A 125 3500 3463 1152 7 1152 $<$1 1152 $<$1 1152 $<$1 1152 $<$1
C250.9_A 250 14017 13967 1236 343 1236 3 1236 3 1236 $<$1 1236 $<$1
gen200_p0.9_55_A 200 8908 9002 1151 113 1151 $<$1 1151 3 1151 $<$1 1151 $<$1
gen400_p0.9_75_A 400 35823 35997 1335 8467 1335 69 1335 2 1335 $<$1 1335 $<$1
hamming8-4_A 256 10329 10535 2155 10800 2155 166 2155 1 2155 $<$1 2155 $<$1
keller4_A 171 4738 4697 1759 353 1759 5 1759 $<$1 1759 $<$1 1759 $<$1
MANN_a27_A 378 35345 35206 1226 1950 1226 10 1226 12 1226 $<$1 1226 $<$1
p_hat300-1_A 300 5505 5428 4740 10800 5637 10800 5637 $<$1 5637 $<$1 5637 $<$1
p_hat300-2_A 300 11051 10877 3342 10800 3750 10800 3943 $<$1 3943 $<$1 3943 $<$1
p_hat300-3_A 300 16820 16570 1658 7243 1658 157 1658 $<$1 1658 $<$1 1658 $<$1
brock200_2_B 200 4916 4960 984 10800 984 757 984 $<$1 984 $<$1 984 $<$1
brock400_2_B 400 29711 30075 765 10800 780 1385 780 10 780 $<$1 780 $<$1
C125.9_B 125 3500 3463 552 7 552 $<$1 552 $<$1 552 $<$1 552 $<$1
C250.9_B 250 14017 13967 586 391 586 3 586 4 586 $<$1 586 $<$1
gen200_p0.9_55_B 200 8908 9002 551 121 551 $<$1 551 1 551 $<$1 551 $<$1
gen400_p0.9_75_B 400 35823 35997 635 9395 635 72 635 $<$1 635 $<$1 635 $<$1
hamming8-4_B 256 10329 10535 1055 10800 1055 163 1055 $<$1 1055 $<$1 1055 $<$1
keller4_B 171 4738 4697 859 367 859 5 859 $<$1 859 $<$1 859 $<$1
MANN_a27_B 378 35345 35206 576 2161 576 11 576 13 576 $<$1 576 $<$1
p_hat300-1_B 300 5505 5428 2340 10800 2325 10800 2787 $<$1 2787 $<$1 2787 $<$1
p_hat300-2_B 300 11051 10877 1642 10800 1850 10800 1943 $<$1 1943 $<$1 1943 $<$1
p_hat300-3_B 300 16820 16570 808 7790 808 161 808 $<$1 808 $<$1 808 $<$1
brock200_2_C 200 4916 4960 459 10800 459 934 459 $<$1 459 $<$1 459 $<$1
brock400_2_C 400 29711 30075 341 10800 355 1711 355 12 355 $<$1 355 $<$1
C125.9_C 125 3500 3463 252 8 252 $<$1 252 $<$1 252 $<$1 252 $<$1
C250.9_C 250 14017 13967 261 501 261 3 261 3 261 $<$1 261 $<$1
gen200_p0.9_55_C 200 8908 9002 251 146 251 1 251 2 251 $<$1 251 $<$1
gen400_p0.9_75_C 400 35823 35997 285 10800 285 83 285 2 285 $<$1 285 $<$1
hamming8-4_C 256 10329 10535 505 10800 505 170 505 $<$1 505 $<$1 505 $<$1
keller4_C 171 4738 4697 409 389 409 5 409 $<$1 409 $<$1 409 $<$1
MANN_a27_C 378 35345 35206 251 2666 251 13 251 10 251 $<$1 251 $<$1
p_hat300-1_C 300 5505 5428 1138 10800 1125 10800 1362 $<$1 1362 $<$1 1362 $<$1
p_hat300-2_C 300 11051 10877 788 10800 900 10800 943 $<$1 943 $<$1 943 $<$1
p_hat300-3_C 300 16820 16570 383 9014 383 178 383 $<$1 383 $<$1 383 $<$1
@@
brock200_2_A 200 2438 7438 1615 10800 1851 10800 1885 $<$1 1885 $<$1 1885 $<$1
brock400_2_A 400 14751 45035 1403 10800 1667 10800 1728 5 1728 $<$1 1728 $<$1
C125.9_A 125 1733 5230 1023 247 1023 4 1023 $<$1 1023 $<$1 1023 $<$1
C250.9_A 250 7073 20911 1200 10800 1236 5143 1236 2 1236 $<$1 1236 $<$1
gen200_p0.9_55_A 200 4425 13485 1186 10800 1206 601 1206 $<$1 1206 $<$1 1206 $<$1
gen400_p0.9_75_A 400 18063 53757 1155 10800 1490 10800 1490 $<$1 1490 $<$1 1490 $<$1
hamming8-4_A 256 5173 15691 1549 10800 1723 10800 1759 $<$1 1759 $<$1 1759 $<$1
keller4_A 171 2400 7035 1400 10800 1434 1573 1434 $<$1 1434 $<$1 1434 $<$1
MANN_a27_A 378 17580 52971 1313 10800 1323 10800 1323 $<$1 1323 $<$1 1323 $<$1
p_hat300-1_A 300 2734 8199 2504 10800 3932 10800 4164 $<$1 4164 $<$1 4164 $<$1
p_hat300-2_A 300 5603 16325 2120 10800 2902 10800 2990 1 2990 $<$1 2990 $<$1
p_hat300-3_A 300 8388 25002 1350 10800 1561 10800 1564 $<$1 1564 $<$1 1564 $<$1
brock200_2_B 200 2438 7438 1416 10800 1602 10800 1641 $<$1 1641 $<$1 1641 $<$1
brock400_2_B 400 14751 45035 1129 10800 1341 10800 1386 2 1386 $<$1 1386 $<$1
C125.9_B 125 1733 5230 856 395 856 8 856 $<$1 856 $<$1 856 $<$1
C250.9_B 250 7073 20911 960 10800 1001 10800 1001 1 1001 $<$1 1001 $<$1
gen200_p0.9_55_B 200 4425 13485 961 10800 983 1585 983 $<$1 983 $<$1 983 $<$1
gen400_p0.9_75_B 400 18063 53757 920 10800 1105 10800 1120 $<$1 1120 $<$1 1120 $<$1
hamming8-4_B 256 5173 15691 1356 10800 1486 10800 1579 $<$1 1579 $<$1 1579 $<$1
keller4_B 171 2400 7035 1230 10800 1268 4468 1268 $<$1 1268 $<$1 1268 $<$1
MANN_a27_B 378 17580 52971 922 10800 1011 10800 1021 $<$1 1021 $<$1 1021 $<$1
p_hat300-1_B 300 2734 8199 2306 10800 3547 10800 3886 $<$1 3886 $<$1 3886 $<$1
p_hat300-2_B 300 5603 16325 1762 10800 2588 10800 2782 $<$1 2782 $<$1 2782 $<$1
p_hat300-3_B 300 8388 25002 1144 10800 1209 10800 1299 $<$1 1299 $<$1 1299 $<$1
brock200_2_C 200 2438 7438 1158 10800 1303 10800 1321 $<$1 1321 $<$1 1321 $<$1
brock400_2_C 400 14751 45035 793 10800 961 10800 1033 $<$1 1033 $<$1 1033 $<$1
C125.9_C 125 1733 5230 644 10800 644 45 644 $<$1 644 $<$1 644 $<$1
C250.9_C 250 7073 20911 688 10800 734 10800 734 1 734 $<$1 734 $<$1
gen200_p0.9_55_C 200 4425 13485 707 10800 727 9879 727 $<$1 727 $<$1 727 $<$1
gen400_p0.9_75_C 400 18063 53757 613 10800 756 10800 772 1 772 $<$1 772 $<$1
hamming8-4_C 256 5173 15691 1120 10800 1248 10800 1378 $<$1 1378 $<$1 1378 $<$1
keller4_C 171 2400 7035 1060 10800 1109 10800 1109 $<$1 1109 $<$1 1109 $<$1
MANN_a27_C 378 17580 52971 583 10800 648 10800 651 2 651 $<$1 651 $<$1
p_hat300-1_C 300 2734 8199 1866 10800 3191 10800 3480 $<$1 3480 $<$1 3480 $<$1
p_hat300-2_C 300 5603 16325 1511 10800 2249 10800 2473 $<$1 2473 $<$1 2473 $<$1
p_hat300-3_C 300 8388 25002 886 10800 919 10800 1004 $<$1 1004 $<$1 1004 $<$1

### v2

@@ B Comparative Details on 216 Instances of SET-1
brock200_2_A 200 4,916 4,960 1,298 8,390 1,298 93 1,298 <<1 1,298 <<1 1,298 <<1
brock400_2_A 400 29,711 30,075 1,103 10,800 1,123 355 1,123 <<1 1,123 <<1 1,123 <<1
C125.9_A 125 3,500 3,463 627 5 627 <<1 627 <<1 627 <<1 627 <<1
C250.9_A 250 14,017 13,967 817 266 817 1 817 <<1 817 <<1 817 <<1
gen200_p0.9_55_A 200 8,908 9,002 785 68 785 <<1 785 <<1 785 <<1 785 <<1
gen400_p0.9_75_A 400 35,823 35,997 895 5,699 895 29 895 <<1 895 <<1 895 <<1
hamming8-4_A 256 10,329 10,535 1,301 7,608 1,301 49 1,301 <<1 1,301 <<1 1,301 <<1
keller4_A 171 4,738 4,697 1,118 153 1,118 2 1,118 <<1 1,118 <<1 1,118 <<1
MANN_a27_A 378 35,345 35,206 812 1,217 812 6 812 <<1 812 <<1 812 <<1
p_hat300-1_A 300 5,505 5,428 2,568 10,800 3,026 10,800 3,129 <<1 3,129 <<1 3,129 <<1
p_hat300-2_A 300 11,051 10,877 2,063 10,800 2,477 10,800 2,477 <<1 2,477 <<1 2,477 <<1
p_hat300-3_A 300 16,820 16,570 1,029 3,748 1,029 38 1,029 <<1 1,029 <<1 1,029 <<1
brock200_2_B 200 4,916 4,960 1,224 9,589 1,224 125 1,224 <<1 1,224 <<1 1,224 <<1
brock400_2_B 400 29,711 30,075 1,010 10,800 1,035 445 1,035 <<1 1,035 <<1 1,035 <<1
C125.9_B 125 3,500 3,463 582 5 582 <<1 582 <<1 582 <<1 582 <<1
C250.9_B 250 14,017 13,967 744 292 744 2 744 <<1 744 <<1 744 <<1
gen200_p0.9_55_B 200 8,908 9,002 716 72 716 <<1 716 <<1 716 <<1 716 <<1
gen400_p0.9_75_B 400 35,823 35,997 805 6,602 805 34 805 <<1 805 <<1 805 <<1
hamming8-4_B 256 10,329 10,535 1,255 7,780 1,255 62 1,255 <<1 1,255 <<1 1,255 <<1
keller4_B 171 4,738 4,697 1,094 156 1,094 2 1,094 <<1 1,094 <<1 1,094 <<1
MANN_a27_B 378 35,345 35,206 707 1,417 707 7 707 1 707 <<1 707 <<1
p_hat300-1_B 300 5,505 5,428 2,492 10,800 2,950 10,800 3,023 <<1 3,023 <<1 3,023 <<1
p_hat300-2_B 300 11,051 10,877 1,979 10,800 2,405 10,800 2,405 <<1 2,405 <<1 2,405 <<1
p_hat300-3_B 300 16,820 16,570 967 4,067 967 46 967 <<1 967 <<1 967 <<1
brock200_2_C 200 4,916 4,960 1,101 10,800 1,101 204 1,101 <<1 1,101 <<1 1,101 <<1
brock400_2_C 400 29,711 30,075 861 10,800 892 691 892 <<1 892 <<1 892 <<1
C125.9_C 125 3,500 3,463 506 6 506 <<1 506 <<1 506 <<1 506 <<1
C250.9_C 250 14,017 13,967 623 334 623 2 623 <<1 623 <<1 623 <<1
gen200_p0.9_55_C 200 8,908 9,002 597 91 597 <<1 597 <<1 597 <<1 597 <<1
gen400_p0.9_75_C 400 35,823 35,997 651 7,858 651 50 651 <<1 651 <<1 651 <<1
hamming8-4_C 256 10,329 10,535 1,184 7,447 1,184 85 1,184 <<1 1,184 <<1 1,184 <<1
keller4_C 171 4,738 4,697 1,049 145 1,049 3 1,049 <<1 1,049 <<1 1,049 <<1
MANN_a27_C 378 35,345 35,206 552 1,849 552 10 552 1 552 <<1 552 <<1
p_hat300-1_C 300 5,505 5,428 2,303 10,800 2,818 10,800 2,897 <<1 2,897 <<1 2,897 <<1
p_hat300-2_C 300 11,051 10,877 1,852 10,800 2,263 10,800 2,263 <<1 2,263 <<1 2,263 <<1
p_hat300-3_C 300 16,820 16,570 851 4,709 851 69 851 <<1 851 <<1 851 <<1
@@
brock200_2_A 200 4,916 4,960 2,034 10,800 2,034 677 2,034 <<1 2,034 <<1 2,034 <<1
brock400_2_A 400 29,711 30,075 1,628 10,800 1,630 1,323 1,630 11 1,630 <<1 1,630 <<1
C125.9_A 125 3,500 3,463 1,152 7 1,152 <<1 1,152 <<1 1,152 <<1 1,152 <<1
C250.9_A 250 14,017 13,967 1,236 343 1,236 3 1,236 3 1,236 <<1 1,236 <<1
gen200_p0.9_55_A 200 8,908 9,002 1,151 113 1,151 <<1 1,151 3 1,151 <<1 1,151 <<1
gen400_p0.9_75_A 400 35,823 35,997 1,335 8,467 1,335 69 1,335 2 1,335 <<1 1,335 <<1
hamming8-4_A 256 10,329 10,535 2,155 10,800 2,155 166 2,155 1 2,155 <<1 2,155 <<1
keller4_A 171 4,738 4,697 1,759 353 1,759 5 1,759 <<1 1,759 <<1 1,759 <<1
MANN_a27_A 378 35,345 35,206 1,226 1,950 1,226 10 1,226 12 1,226 <<1 1,226 <<1
p_hat300-1_A 300 5,505 5,428 4,740 10,800 5,637 10,800 5,637 <<1 5,637 <<1 5,637 <<1
p_hat300-2_A 300 11,051 10,877 3,342 10,800 3,750 10,800 3,943 <<1 3,943 <<1 3,943 <<1
p_hat300-3_A 300 16,820 16,570 1,658 7,243 1,658 157 1,658 <<1 1,658 <<1 1,658 <<1
brock200_2_B 200 4,916 4,960 984 10,800 984 757 984 <<1 984 <<1 984 <<1
brock400_2_B 400 29,711 30,075 765 10,800 780 1,385 780 10 780 <<1 780 <<1
C125.9_B 125 3,500 3,463 552 7 552 <<1 552 <<1 552 <<1 552 <<1
C250.9_B 250 14,017 13,967 586 391 586 3 586 4 586 <<1 586 <<1
gen200_p0.9_55_B 200 8,908 9,002 551 121 551 <<1 551 1 551 <<1 551 <<1
gen400_p0.9_75_B 400 35,823 35,997 635 9,395 635 72 635 <<1 635 <<1 635 <<1
hamming8-4_B 256 10,329 10,535 1,055 10,800 1,055 163 1,055 <<1 1,055 <<1 1,055 <<1
keller4_B 171 4,738 4,697 859 367 859 5 859 <<1 859 <<1 859 <<1
MANN_a27_B 378 35,345 35,206 576 2,161 576 11 576 13 576 <<1 576 <<1
p_hat300-1_B 300 5,505 5,428 2,340 10,800 2,325 10,800 2,787 <<1 2,787 <<1 2,787 <<1
p_hat300-2_B 300 11,051 10,877 1,642 10,800 1,850 10,800 1,943 <<1 1,943 <<1 1,943 <<1
p_hat300-3_B 300 16,820 16,570 808 7,790 808 161 808 <<1 808 <<1 808 <<1
brock200_2_C 200 4,916 4,960 459 10,800 459 934 459 <<1 459 <<1 459 <<1
brock400_2_C 400 29,711 30,075 341 10,800 355 1,711 355 12 355 <<1 355 <<1
C125.9_C 125 3,500 3,463 252 8 252 <<1 252 <<1 252 <<1 252 <<1
C250.9_C 250 14,017 13,967 261 501 261 3 261 3 261 <<1 261 <<1
gen200_p0.9_55_C 200 8,908 9,002 251 146 251 1 251 2 251 <<1 251 <<1
gen400_p0.9_75_C 400 35,823 35,997 285 10,800 285 83 285 2 285 <<1 285 <<1
hamming8-4_C 256 10,329 10,535 505 10,800 505 170 505 <<1 505 <<1 505 <<1
keller4_C 171 4,738 4,697 409 389 409 5 409 <<1 409 <<1 409 <<1
MANN_a27_C 378 35,345 35,206 251 2,666 251 13 251 10 251 <<1 251 <<1
p_hat300-1_C 300 5,505 5,428 1,138 10,800 1,125 10,800 1,362 <<1 1,362 <<1 1,362 <<1
p_hat300-2_C 300 11,051 10,877 788 10,800 900 10,800 943 <<1 943 <<1 943 <<1
p_hat300-3_C 300 16,820 16,570 383 9,014 383 178 383 <<1 383 <<1 383 <<1
@@
brock200_2_A 200 2,438 7,438 1,615 10,800 1,851 10,800 1,885 <<1 1,885 <<1 1,885 <<1
brock400_2_A 400 14,751 45,035 1,403 10,800 1,667 10,800 1,728 5 1,728 <<1 1,728 <<1
C125.9_A 125 1,733 5,230 1,023 247 1,023 4 1,023 <<1 1,023 <<1 1,023 <<1
C250.9_A 250 7,073 20,911 1,200 10,800 1,236 5,143 1,236 2 1,236 <<1 1,236 <<1
gen200_p0.9_55_A 200 4,425 13,485 1,186 10,800 1,206 601 1,206 <<1 1,206 <<1 1,206 <<1
gen400_p0.9_75_A 400 18,063 53,757 1,155 10,800 1,490 10,800 1,490 <<1 1,490 <<1 1490 <<1
hamming8-4_A 256 5,173 15,691 1,549 10,800 1,723 10,800 1,759 <<1 1,759 <<1 1,759 <<1
keller4_A 171 2,400 7,035 1,400 10,800 1,434 1573 1,434 <<1 1,434 <<1 1,434 <<1
MANN_a27_A 378 17,580 52,971 1,313 10,800 1,323 10,800 1,323 <<1 1,323 <<1 1,323 <<1
p_hat300-1_A 300 2,734 8,199 2,504 10,800 3,932 10,800 4,164 <<1 4,164 <<1 4,164 <<1
p_hat300-2_A 300 5,603 16,325 2,120 10,800 2,902 10,800 2,990 1 2,990 <<1 2,990 <<1
p_hat300-3_A 300 8,388 25,002 1,350 10,800 1561 10,800 1,564 <<1 1,564 <<1 1,564 <<1
brock200_2_B 200 2,438 7,438 1,416 10,800 1,602 10,800 1,641 <<1 1,641 <<1 1,641 <<1
brock400_2_B 400 14,751 45,035 1,129 10,800 1,341 10,800 1,386 2 1,386 <<1 1,386 <<1
C125.9_B 125 1,733 5,230 856 395 856 8 856 <<1 856 <<1 856 <<1
C250.9_B 250 7,073 20,911 960 10,800 1,001 10,800 1,001 1 1,001 <<1 1,001 <<1
gen200_p0.9_55_B 200 4,425 13,485 961 10,800 983 1585 983 <<1 983 <<1 983 <<1
gen400_p0.9_75_B 400 18,063 53,757 920 10,800 1,105 10,800 1,120 <<1 1,120 <<1 1,120 <<1
hamming8-4_B 256 5,173 15,691 1,356 10,800 1,486 10,800 1,579 <<1 1,579 <<1 1,579 <<1
keller4_B 171 2,400 7,035 1,230 10,800 1,268 4468 1,268 <<1 1,268 <<1 1,268 <<1
MANN_a27_B 378 17,580 52,971 922 10,800 1,011 10,800 1,021 <<1 1,021 <<1 1,021 <<1
p_hat300-1_B 300 2,734 8,199 2,306 10,800 3,547 10,800 3,886 <<1 3,886 <<1 3,886 <<1
p_hat300-2_B 300 5,603 16,325 1,762 10,800 2,588 10,800 2,782 <<1 2,782 <<1 2,782 <<1
p_hat300-3_B 300 8,388 25,002 1,144 10,800 1,209 10,800 1,299 <<1 1,299 <<1 1,299 <<1
brock200_2_C 200 2,438 7,438 1,158 10,800 1,303 10,800 1,321 <<1 1,321 <<1 1,321 <<1
brock400_2_C 400 14,751 45,035 793 10,800 961 10,800 1,033 <<1 1,033 <<1 1,033 <<1
C125.9_C 125 1,733 5,230 644 10,800 644 45 644 <<1 644 <<1 644 <<1
C250.9_C 250 7,073 20,911 688 10,800 734 10,800 734 1 734 <<1 734 <<1
gen200_p0.9_55_C 200 4,425 13,485 707 10,800 727 9,879 727 <<1 727 <<1 727 <<1
gen400_p0.9_75_C 400 18,063 53,757 613 10,800 756 10,800 772 1 772 <<1 772 <<1
hamming8-4_C 256 5,173 15,691 1,120 10,800 1,248 10,800 1,378 <<1 1,378 <<1 1,378 <<1
keller4_C 171 2,400 7,035 1,060 10,800 1,109 10,800 1,109 <<1 1,109 <<1 1,109 <<1
MANN_a27_C 378 17,580 52,971 583 10,800 648 10,800 651 2 651 <<1 651 <<1
p_hat300-1_C 300 2,734 8,199 1,866 10,800 3,191 10,800 3,480 <<1 3,480 <<1 3,480 <<1
p_hat300-2_C 300 5,603 16,325 1,511 10,800 2,249 10,800 2,473 <<1 2,473 <<1 2,473 <<1
p_hat300-3_C 300 8,388 25,002 886 10,800 919 10,800 1,004 <<1 1,004 <<1 1,004 <<1